OpenStudy (anonymous):

Simplify (5a7b)^2. I got \(25a^{14}b^2\)

OpenStudy (anonymous):

@mathslover @nincompoop

OpenStudy (anonymous):

@Luigi0210

OpenStudy (kendricklamar2014):

\[25a^14b^2~is~Correct\]

OpenStudy (anonymous):

Haha thnx :D

OpenStudy (luigi0210):

That would be right. \(\Large (x^a)^b=x^{a*b}\) Like Ken said xD

OpenStudy (anonymous):

I have another question. \(x^2x^4=x^{2+4}=x^6\) is the product of power right

OpenStudy (anonymous):

product of powers

OpenStudy (kendricklamar2014):

Yes, you are correct. You add up the powers.